# A Novel Metric for Measuring the Robustness of Large Language Models in Non-adversarial Scenarios

### v1

@@ Abstract
We evaluate the robustness of several large language models on multiple datasets. Robustness here refers to the relative insensitivity of the model’s answers to meaning-preserving variants of their input. Benchmark datasets are constructed by introducing naturally-occurring, non-malicious perturbations, or by generating semantically equivalent paraphrases of input questions or statements. We further propose a novel metric for assessing a model robustness, and demonstrate its benefits in the non-adversarial scenario by empirical evaluation of several models on the created datasets.
@@ 1 Introduction
Benchmarking the robustness of a model to variations in its input typically involves measuring the degree of performance decrease in perturbed instance set, compared to the original example. Assessing the resilience of LLMs to adversarial attacks, the main metric that has been put forward in prior studies is performance drop rate (PDR), which is the fractional decrease in the average perturbed instances’ score, relative to the original example (Zhu et al., 2023). As discussed in Section 3.1, PDR has two main drawbacks: First, since it measures fractional decrease, it is inherently an asymmetric function of its inputs; a fixed increase in performance after perturbation, receives a larger magnitude PDR than the reversed decrease in performance. Second, since fractional change from 0 is undefined, the PDR is undefined in the specific case when the original score was zero but the average over perturbed instance set scores higher, as in the example in Table 1; thus, instances with undefined PDR are ignored when evaluating average PDR on a dataset (e.g., Liang et al. (2023); Zhu et al. (2023)), which can bias aggregates.
@@ 2.2 Expanding Datasets with Perturbations
Simple non-semantic perturbations, such as upper-, lower- or proper-casing of certain words, removing punctuation, "butter-finger" typos (misspelling by replacing a randomly-selected letter with one of the adjacent ones on a keyboard), character swap, redundant white-spacing. A sentence variant can include one or multiple interventions from this set, as illustrated in examples (2)-(4) in Table 1.
@@
Table 2 summarizes the datasets statistics before and after expansion, and perturbation types used. Let $\mathcal{D}$ denote an unperturbed test dataset, consisting of $n$ unique instances $x_{1},\dots,x_{n}$ (e.g., questions to be answered). The expanded dataset $\mathcal{D}^{\prime}$ consists of each original instance $x_{i}$ (now denoted by $x_{i}^{o}$), as well as the set of its $m(i){\geq}1$ perturbations, denoted by $(x_{i}^{1},\dots,x_{i}^{m(i)})$.
dataset expansion type orig. size final size
PopQA S, P 14,268 240,246
BoolQ S, P, D 3,271 29,394
SIGA S, P 2,122 20,833
Table 2: Datasets statistics before and after expansion. Only test set portion was considered where applicable.
@@ 3 Quantifying Model Robustness
Table 3: Mean accuracy on the original datasets, mean accuracy on perturbed variants (slightly lower). The high difference between the directional $\tilde{\textrm{H}}$ and undirectional $\textrm{A}\tilde{\textrm{H}}$ is suggestive of both increase and decrease in models’ performance on original, compared to perturbed examples. Results for which the $\tilde{\textrm{H}}$ and $\textrm{A}\tilde{\textrm{H}}$ 95% confidence intervals (CIs, see Appendix A.2.) do not contain 0 are marked with "*", indicating the significance of the finding. Notably, significant $\tilde{\textrm{H}}$ and $\textrm{A}\tilde{\textrm{H}}$ values may still indicate a very small effect size (see Appendix A.1); values reflecting a non-negligible change are marked with gray background. The best result in a column is boldfaced.
Consider $x_{i}^{o}$ and $(x_{i}^{1},\dots,x_{i}^{m(i)})$, the $i^{\text{th}}$ original input and its $m(i)$ perturbations, in the perturbed dataset $\mathcal{D}^{\prime}$ (see Section 2.2). Let $score_{i}^{o}$ and $(score_{i}^{1},\dots,score_{i}^{m(i)})$ be binary-valued scores (accuracy or string containment match) of the ground truth reference to the model’s predicted value (e.g., generated answer) on the original ("o") and perturbed instances, respectively. We compute the average score on the perturbed ("p") instance set for input $x_{i}^{o}$ in $\mathcal{D}^{\prime}$ by $score_{i}^{p}=\frac{1}{m(i)}\sum_{j=1}^{m(i)}score_{i}^{j}$. We consider a model’s performance on a dataset as being robust if the performance tends to be insensitive to perturbations; that is, the two scores—$score_{i}^{o}$ and $score_{i}^{p}$—tend to be close to to each other, across all input instances $i$ in $\mathcal{D}$.
3.1 Performance Drop Rate (PDR)
3.2 Cohen’s $h$ Effect Size
Cohen’s $h$ Cohen (1988) statistical effect size is commonly used for measuring the difference in two proportions in empirical research setup (see Appendix A.1 for background), and is defined as
Cohen’s $h$ effect size takes values in the range $[-\pi,\pi]$, where $h{>}0$ indicates performance improvement relative to $score_{i}^{o}$. This metric has several important characteristics: (1) Unlike PDR, it is a symmetric function, and is defined for all pairs of $score$ values within the [0, 1] range. (2) It has rule-of-thumb thresholds [^4] See Table 4 in Appendix  A.1. of a statistically-significant difference in sample proportions, which are based on its statistical properties. For better interpretability, we define a normalized version of Cohen’s $h$, defined as $\tilde{\textrm{H}}=\textrm{H}/\pi$, which takes values within the [-1, 1] range. Consequently, the significance thresholds are adjusted for this normalized version, each divided by $\pi$. Figure 2 in the Appendix A.1 shows that PDR and $\tilde{\textrm{H}}$ correlate very well (Pearson’s $r{\approx}0.995$ when $score_{i}^{o}{=}1.0$), supporting this novel application of the metric to the scenario of tasks with binary-valued evaluation outcome.
4 Benchmarking Model Robustness
4.1 Experimental Setup
4.2 Experimental Results
Table 3 shows per-dataset performance by model. Original performance (average on examples in the raw dataset – mean($score_{i}^{o}$)) is reported, as well as the average over mean perturbed sets – mean($score_{i}^{p}$). A slightly inferior performance on the perturbed instances compared to original is reflected by the negative Cohen’s $h$ effect size ($\tilde{\textrm{H}}$). Notably, the absolute value effect ($\textrm{A}\tilde{\textrm{H}}$) differs considerably from the directional $\tilde{\textrm{H}}$ values, indicating that the LLMs’ predictions varied in both directions (better or worse) compared to the original instance performance – a finding largely supportive of our assumption that naturally-occurring, non-malicious perturbations may have either positive or negative effect on a model’s accuracy. While only a single observed $\tilde{\textrm{H}}$ directional effect size is considered non-negligible (grey background), most absolute values constitute a "small" change, and virtually all effect sizes show significant, indicating small, yet systematic, and reliably detected change.
@@ 4.2 Experimental Results
Figure 1 illustrates average model accuracy on the original datasets vs their mean undirectional robustness ($\textrm{A}\tilde{\textrm{H}}$). Evidently, best performing models (Llama3-Instruct (70B) and Mistral-Instruct (8$\times$7B)) exhibit more moderate robustness, compared to the most robust Flan-T5-XXL, that shows slightly inferior average performance. Granite-Instruct (13B) is one of the most robust models, while performing worse on average.
@@ A.1 Effect Sizes and Cohen’s $h$
Effect size metrics are measures of the size of a statistical phenomenon; common examples are Pearson correlation and odds ratios. An effect size metric measures the aspect of interest (e.g., the difference in means between two sample sizes) in a way that is independent of the sample sizes. This is in contrast to the p-value of a hypothesis test statistic (e.g., a two-sample test) which, for fixed values of the sample means and variances, becomes more significant when the sample sizes increase Sullivan and Feinn (2012); this quality makes p-values vulnerable to manipulation ("p-hacking"). Effect size metrics are often used to ensure that a hypothesis test has enough statistical power (complement of the Type-II or false negative error probability) given the sample size(s). This insensitivity to the sample size in the effect size value means that they can be used to measure the significance of an effect both in cases of small sample sizes (e.g., in our case when we have one original instance and a small number of perturbations), and thus may be better than, say, a two-sample p-value, for assessing robustness. Cohen’s $h$ has a particular advantage in that it is defined if there is no sample variation (e.g., if $score\in\{0,1\}$), which causes p-values and some other effect sizes to be undefined.
Cohen’s $h$—and thus $\tilde{\textrm{H}}$—changes non-linearly with changes in $|score_{i}^{o}{-}score_{i}^{p}|$. In contrast, PDR changes linearly when $score_{i}^{o}$ is fixed. Considering the specific case when the original instance accuracy $score_{i}^{o}$ is perfect (1.0), $\tilde{\textrm{H}}$ and "reverse PDR" (${=}-1{\times}\textrm{PDR}$) both metrics take values in the [-1, 0] range and are highly correlated, as shown in Figure 2, suggesting that in this case ($score_{i}^{o}{=}1.0$), Cohen’s $h$ constitutes an intuitive and easily-interpretable alternative to PDR.
@@
We note that if the instance scores $score_{i}^{j}$ are continuous-valued rather than binary, an alternative effect size metric such as Cohen’s $d$ Cohen (1988) or Hedges’ $g$ Hedges (1981) for comparing sample means can be used in a way, similar to Cohen’s $h$. However, these effect sizes are undefined or infinite when the within-sample variance is zero. We leave further investigation for future work.
@@ A.2 Bootstrapped Confidence Intervals
All metrics and instance scoring functions are implemented in the open-source repository unitxt Bandel et al. (2024). A group here consists of an original instance and its $m(i)$ perturbations (see Section 2.2). A given metric $f$ (e.g., mean score, PDR, $\tilde{\textrm{H}}$) produces a group-level score $s_{i}{=}f\left(score_{i}^{o},score_{i}^{1},\dots,score_{i}^{m(i)}\right)$. Thus, we have instance-group scores $(s_{1},\dots,s_{n})$ on $D^{\prime}$. Statistical analysis of the metric is done by constructing 95% bootstrapped confidence intervals on the $s_{i}$ scores, discarding any undefined values, rather than resampling the instance scores, reforming the groups, and calculating the group scores; the latter option could result in duplicated perturbed instances or incomplete groups if $score_{i}^{o}$ is missing, which can make $s_{i}$ undefined. We conduct group-score resampling because we analyze the typical average robustness across original instances, and thus the unit of analysis is the original instance together with its perturbations, as reflected in $s_{i}$.

### v2

@@ Abstract
We evaluate the robustness of several large language models on multiple datasets. Robustness here refers to the relative insensitivity of the model’s answers to meaning-preserving variants of their input. Benchmark datasets are constructed by introducing naturally-occurring, non-malicious perturbations, or by generating semantically equivalent paraphrases of input questions or statements. We further propose a novel metric for assessing a model robustness, and demonstrate its benefits in the non-adversarial scenario by empirical evaluation of several models on the created datasets.  [^1] All our datasets are available on HuggingFace.
@@ 1 Introduction
Benchmarking the robustness of a model to variations in its input typically involves measuring the degree of performance decrease in the perturbed instance set, compared to the original example. For assessing the resilience of LLMs to adversarial attacks, the main metric that has been put forward in prior studies is performance drop rate (PDR), which is the fractional decrease in the average perturbed instances’ score, relative to the original example (Zhu et al., 2023). As discussed in Section 3.1, PDR has two main drawbacks: First, since it measures fractional decrease, it is inherently an asymmetric function of its inputs; a fixed increase in performance after perturbation receives a larger magnitude PDR than the reversed decrease in performance. Second, since fractional change from 0 is undefined, the PDR is undefined in the specific case when the original score was zero but the average over perturbed instance set scores higher, as in the example in Table 1; thus, instances with undefined PDR are ignored when evaluating average PDR on a dataset (e.g., Liang et al. (2023); Zhu et al. (2023)), which can bias aggregates.
@@
Much effort has been invested recently in leaderboards for multi-faceted evaluation of foundation models (e.g., Liang et al. (2023) [^3] https://crfm.stanford.edu/helm/). A broader impact of this study lies in the adoption of the proposed metric for benchmarking the robustness of LLMs in non-adversarial scenarios.
@@ 2.2 Expanding Datasets with Perturbations
Simple non-semantic perturbations, such as upper-, lower- or proper-casing of certain words, removing punctuation, "butter-finger" typos (misspelling by replacing a randomly-selected letter with one of the adjacent ones on a keyboard), character swap, or redundant white-spacing. A sentence variant can include one or multiple interventions from this set, as illustrated in examples (2)-(4) in Table 1.
@@
Table 2 summarizes the datasets statistics before and after expansion, and perturbation types used. Let $\mathcal{D}$ denote an unperturbed test dataset, consisting of $n$ unique instances $x_{1},\dots,x_{n}$ (e.g., questions to be answered). The expanded dataset $\mathcal{D}^{\prime}$ consists of each original instance $x_{i}$ (now denoted by $x_{i}^{o}$), as well as the set of its $m(i){\geq}1$ perturbations, denoted by $(x_{i}^{1},\dots,x_{i}^{m(i)})$. Our expanded datasets are available on HuggingFace.
dataset original S P D final
PopQA 14.2K 85.6K 104.3K 204.2K
BoolQ 3.2K 9.8K 9.7K 6.5K 29.3K
SIGA 2.1K 12.6K 6.1K 20.8K
Table 2: Datasets size before and after expansion, by perturbation type: superficial (S), paraphrase (P) and distraction (D). Only test set portion of the dataset was considered for experiments where applicable.
@@ 3 Quantifying Model Robustness
Consider $x_{i}^{o}$ and $(x_{i}^{1},\dots,x_{i}^{m(i)})$, the $i^{\text{th}}$ original input and its $m(i)$ perturbations, in the perturbed dataset $\mathcal{D}^{\prime}$ (see Section 2.2). Given a scoring metric $score\in[0,1]$, let $score_{i}^{o}$ and $(score_{i}^{1},\dots,score_{i}^{m(i)})$ be scores of the model’s predicted value (e.g., generated answer) on the original ("o") and perturbed instances, respectively, versus the ground truth reference. In our case, $score\in\{0,1\}$ are binary-valued accuracy or string containment match metrics. We compute the average score on the perturbed ("p") instance set for input $x_{i}^{o}$ in $\mathcal{D}^{\prime}$ by $score_{i}^{p}=\frac{1}{m(i)}\sum_{j=1}^{m(i)}score_{i}^{j}$. We consider a model’s performance on a dataset as being robust if the performance tends to be insensitive to perturbations; that is, the two scores—$score_{i}^{o}$ and $score_{i}^{p}$—tend to be close to to each other, across all input instances $i$ in $\mathcal{D}$.
3.1 Performance Drop Rate (PDR)
3.2 Cohen’s $h$ Effect Size
Cohen’s $h$ Cohen (1988) statistical effect size is commonly used for measuring the difference in two proportions in empirical research (see Appendix A.1 for background), and is defined as
Cohen’s $h$ effect size takes values in the range $[-\pi,\pi]$, where $h{>}0$ indicates performance improvement relative to $score_{i}^{o}$. This metric has several important characteristics: (1) Unlike PDR, it is a symmetric function, and is defined for all pairs of $score$ values within the [0, 1] range. (2) It has rule-of-thumb thresholds [^6] See Table 5 in Appendix A.1. of what values constitute small, medium, etc. differences in sample proportions, which are based on its statistical properties. For better interpretability, we define a normalized version of Cohen’s $h$, defined as $\tilde{\textrm{H}}=\textrm{H}/\pi$, which takes values within the [-1, 1] range. Consequently, the effect size thresholds are adjusted for this normalized version, each divided by $\pi$.
Figure 1 shows that PDR and $\tilde{\textrm{H}}$ correlate very well (Pearson’s $r{\approx}0.995$ when $score_{i}^{o}{=}1.0$), supporting this novel application of the metric to the scenario of tasks with binary-valued evaluation outcome. We note, however, that the application of the metric is not limited to the scenario with binary evaluation outcome. See Appendix A.1 for a statistical discussion of the $\tilde{\textrm{H}}$ metric.
Taking the example values of $score_{i}^{o}=0.8$ and $score_{i}^{p}=0.1$ used above for PDR (Section 3.1), the corresponding value of $\tilde{\textrm{H}}$ is -0.5, and 0.5 if the direction is reversed (non-normalized $\textrm{H}\approx\pm 1.57$). This counts as a ‘very large’ difference according to the thresholds in Table 5 (see Appendix A.1).
4 Benchmarking Model Robustness
4.1 Experimental Setup
4.2 Experimental Results
Table 3 shows per-dataset performance by model. Original performance (average on examples in the raw dataset – mean($score_{i}^{o}$)) is reported, as well as the average over mean perturbed sets – mean($score_{i}^{p}$). Figure 3 in Appendix A.3 visualizes this table along with the 95% confidence intervals for the metrics, represented by the cell shading in the table. The significance thresholds in Table 5 can tell us how large the observed mean $\tilde{\textrm{H}}$ across examples in a dataset is, and one can decide that, say, a difference that is ‘small’ or greater indicates the LLM is not robust to the perturbations. The 95% interval for $\tilde{\textrm{H}}$ (or $\textrm{A}\tilde{\textrm{H}}$) measure how variable this assessment of robustness is to random sampling; for instance, if the 95% interval is itself also within the bounds of a ‘small’ difference, this lends statistical confidence to the assessment that the LLM is robust. The fact that $\tilde{\textrm{H}}$ has such significance thresholds that can be used in a practical decision gives it a benefit compared to metrics like PDR, which lack them.
A slightly inferior performance on the perturbed instances compared to original is reflected by the negative Cohen’s $h$ effect size ($\tilde{\textrm{H}}$). Notably, the absolute value effect ($\textrm{A}\tilde{\textrm{H}}$) differs considerably from the directional $\tilde{\textrm{H}}$ values, indicating that the LLMs’ predictions varied in both directions (better or worse) compared to the original instance performance – a finding largely supportive of our assumption that naturally-occurring, non-malicious perturbations may have either positive or negative effect on a model’s accuracy. While only a single observed $\tilde{\textrm{H}}$ directional effect size is considered non-negligible (grey background), most absolute values constitute a "small" change, and virtually all effect sizes show significant, indicating small, yet systematic, and reliably detected change.
@@ 4.2 Experimental Results
Table 3: Mean accuracy on the original datasets, mean accuracy on perturbed variants (slightly lower). Confidence intervals (CIs) for are calculated by original or perturbed group-level bootstrapping, as discussed in Appendix A.2. The high difference between the directional $\tilde{\textrm{H}}$ and undirectional $\textrm{A}\tilde{\textrm{H}}$ is suggestive of both increase and decrease in models’ performance on original, compared to perturbed examples. Results for which the $\tilde{\textrm{H}}$ and $\textrm{A}\tilde{\textrm{H}}$ 95% confidence intervals (see Appendix A.2.) do not contain 0 are marked with "*", indicating the significance of the finding. Notably, significant $\tilde{\textrm{H}}$ and $\textrm{A}\tilde{\textrm{H}}$ values may still indicate a very small effect size (see Appendix A.1); values reflecting a non-negligible change are marked with gray background. The best result in a column is boldfaced.
superficial (S) paraphrase (P) distraction (D)
dataset M(orig) M(pert.) $\tilde{\textrm{H}}$ $\textrm{A}\tilde{\textrm{H}}$ M(pert.) $\tilde{\textrm{H}}$ $\textrm{A}\tilde{\textrm{H}}$ M(pert.) $\tilde{\textrm{H}}$ $\textrm{A}\tilde{\textrm{H}}$
PopQA 0.37 0.32 -0.05* 0.12* 0.34 -0.03* 0.15*
BoolQ 0.89 0.88 -0.01* 0.04* 0.85 -0.04* 0.07* 0.88 -0.01 0.05*
SIGA 0.80 0.79 -0.01* 0.06* 0.77 0.00* 0.09*
Table 4: Mean accuracy on the original datasets, mean accuracy on perturbed variants with the most recent Llama3-Instruct(70B) model in this study, with break-down by variant type (superficial (S), paraphrase (P), distraction (D)). Results for which the $\tilde{\textrm{H}}$ and $\textrm{A}\tilde{\textrm{H}}$ 95% confidence intervals (see Appendix A.2.) do not contain 0 are marked with "*", indicating the significance of the finding; values reflecting a non-negligible change are marked with gray background. Notably, paraphrasing the original question results in a more considerable performance drop than introducing superficial (simple) perturbations, across all three datasets.
@@
Model Robustness vs Performance
Figure 2 illustrates average model accuracy on the original datasets vs their mean undirectional robustness ($\textrm{A}\tilde{\textrm{H}}$). Evidently, best performing, recently released models (Llama3-Instruct (70B) and Mistral-Instruct (8$\times$7B)) exhibit more moderate robustness, compared to the most robust Flan-T5-XXL, that shows slightly inferior average performance. Granite-Instruct (13B) is one of the most robust models, while performing worse on average.
Robustness Evaluation by Perturbation Type
We further break down the robustness measurements by individual perturbation types in Table 4 for Llama3-Instruct(70B) – one of the best-performing models on our datasets. Notably, paraphrasing the original question results in a more considerable performance drop than introducing superficial (simple) perturbations, across all three datasets. We attribute the particularly high absolute effect size in the PopQA dataset (0.15) to the fact that paraphrases in this dataset were created manually, aiming at high linguistic diversity while maintaining the original semantics.
@@ A.1 Effect Sizes and Cohen’s $h$
Effect size metrics are measures of the size of a statistical phenomenon; common examples are Pearson correlation and odds ratios. An effect size metric measures the aspect of interest (e.g., the difference in means between two sample sizes) in a way that is independent of the sample sizes. This is in contrast to the p-value of a hypothesis test statistic (e.g., a two-sample test) which, for fixed values of the sample means and variances, becomes more significant when the sample sizes increase Sullivan and Feinn (2012); this quality makes p-values vulnerable to manipulation ("p-hacking"). Effect size metrics are often used to ensure that a hypothesis test has enough statistical power (complement of the Type-II or false negative error probability) given the sample size(s). This insensitivity to the sample size in the effect size value means that they can be used to measure the significance of an effect in cases of small sample sizes (e.g., in our case when we have one original instance and a small number of perturbations), and thus may be better than, say, a two-sample p-value, for assessing robustness. Cohen’s $h$ has a particular advantage in that it is defined even if there is no sample variation (e.g., if $score\in\{0,1\}$), which causes p-values and some other effect sizes to be undefined.
Cohen’s $h$—and thus $\tilde{\textrm{H}}$—changes non-linearly with changes in $|score_{i}^{o}{-}score_{i}^{p}|$. In contrast, PDR changes linearly when $score_{i}^{o}$ is fixed. Considering the specific case when the original instance accuracy $score_{i}^{o}$ is perfect (1.0), both $\tilde{\textrm{H}}$ and "reverse PDR" (${=}-1{\times}\textrm{PDR}$) take values in the [-1, 0] range and are highly correlated, as shown in Figure 1; this high correlation suggests that in this case ($score_{i}^{o}{=}1.0$), Cohen’s $h$ constitutes an intuitive and easily-interpretable alternative to PDR.
@@
We note that if the instance scores $score_{i}^{j}$ are continuous-valued rather than binary, an alternative effect size metric such as Cohen’s $d$ Cohen (1988) or Hedges’ $g$ Hedges (1981) for comparing sample means can be used in a similar way to Cohen’s $h$. However, these effect sizes, unlike Cohen’s $h$, are undefined or infinite when the within-sample variance is zero. We leave further investigation for future work.
@@ A.2 Bootstrapped Confidence Intervals
All metrics and instance scoring functions are implemented in the open-source repository unitxt Bandel et al. (2024). A ‘group’ here consists of an original instance and its $m(i)$ perturbations (see Section 2.2). A given metric $f$ (e.g., mean score, PDR, $\tilde{\textrm{H}}$) produces a group-level score $s_{i}{=}f\left(score_{i}^{o},score_{i}^{1},\dots,score_{i}^{m(i)}\right)$. Thus, if original dataset $\mathcal{D}$ has $n$ instances, we have $n$ instance-group scores $(s_{1},\dots,s_{n})$ on $\mathcal{D}^{\prime}$. Statistical analysis of the metric is done by constructing 95% bootstrapped confidence intervals on the $s_{i}$ scores, discarding any undefined values, rather than resampling the instance scores, reforming the groups, and calculating the group scores; the latter option could result in duplicated perturbed instances or incomplete groups if $score_{i}^{o}$ is missing, which can make $s_{i}$ undefined. We conduct group-score resampling because we analyze the typical average robustness across original instances, and thus the unit of analysis is the original instance together with its perturbations, as reflected in $s_{i}$.